OpenStudy (anonymous):

Write the standard equation for the circle given the following equations x2 + y2 - 8x + 7 = 0 x2 + y2 + 4x - 6y - 3 = 0

OpenStudy (stamp):

Hint: Complete the square.

OpenStudy (stamp):

Start with the first equation, once we have that established we can move on to the second.

OpenStudy (anonymous):

i add the squares together ?

OpenStudy (stamp):

Completing the square example:\[x^2+4x=0\]\[(x+2)^2=x^2+4x+4\]\[(x+2)^2-4=0\]\[(x+2)^2=4\]

OpenStudy (stamp):

Use completing the square in your equation to get \[(x-h)^2+(y-k)^2=r^2\]which is your standard circle equation. Do you see how completing the square can be helpful to simplify these equations into standard circle form?

OpenStudy (anonymous):

yea

OpenStudy (stamp):

Ok so let me know what you get for your first solution.